OpenStudy (anonymous):

Find additional two roots of 7+ square root of 6 and 2- square root of 1... What are 2 additional roots??

OpenStudy (freckles):

is there not more information?

OpenStudy (freckles):

like are we talking about roots of a polynomial with real coefficients

OpenStudy (anonymous):

No it just says to find additional roots and no idea how to get that @freckles

OpenStudy (anonymous):

Yeah polynomials

OpenStudy (dumbcow):

radical roots always come in pairs \[a \pm \sqrt{b}\] so when they give you one root \[7 + \sqrt{6}\] then another root must be: \[7 - \sqrt{6}\]

OpenStudy (anonymous):

So the next one would be 2+ square root of 1?

OpenStudy (dumbcow):

correct

OpenStudy (anonymous):

Thank you!